OpenStudy (anonymous):

If sin(x) = 1/3 and sec(y) = 5/4 , where x and y lie between 0 and pie/2, evaluate the expression cos(x+y).

OpenStudy (cwrw238):

cos(x + y) = cosx cosy - sinx siny cosy = 1/ secy = 4/5 you can find the value of siny and cosx by using the identities siny = sqrt (1 - cos^2y) and cos x = sqrt(1 - sin^2 x)